In a survey of a TriDelt chapter with 50 members, 21 were taking mathematics, 36 were taking English, and 9 were taking both. How many were not taking either of these subjects?

Answer by Edwin McCravy(20060)

In a survey of a TriDelt chapter with 50 members, 21 were taking mathematics, 36 were taking English, and 9 were taking both. How many were not taking either of these subjects?
We draw a rectangle with two overlapping circles, one red and one blue.
 


There are two circles.  The red one contains all the students taking
math.  The blue one contains all the students taking English.  Note
that the red circle and the blue circle overlap.  

Here is the mistake many students make, so DO NOT make this
mistake:

Many students would see the words "21 were taking mathematics,36 were taking English,",
and put all 21 in region "a", and all 36 in region "c", like this:



DO NOT make that error!!! Some, but not all, of the 21 go in
region "a" and the rest of the go in region "b".  Similar,
Some, but not all, of the 36 go in region "d" and the rest 
of the go in region "b".

So we do not begin with either the 21 or the 36. 

 Instead we begin with  "9 were taking both".  These 9 go in
the middle region, b, because those 9 are in both the red
circle and the blue circle.

So we put the 9 in region b:



Now we see "21 were taking mathematics", and of this 21,
9 are taking English too, so we must subtract the 9
that are in the middle region, getting 21-9 = 12, and
we put 12 in region "a":



Now we see "36 were taking English", and of this 36,
9 are taking  too, so we must subtract the 9
that are in the middle region, getting 36-9 = 27, and
we put 27 in region "c":


 
We have one more region, "d".  That is the region
outside of both circles.  They are the ones that are neither
taking mathematics nor taking English.

We have accounted for 12+9+27 = 48.  We are told there are 50
members. So there are 50-48 = 2 that have not been accounted
for.  Those 2 go in the outer region "d":



So the answer to the problem is those 2.

Edwin